what is the equation to solving this kind of problem?
At 1:00 PM Lucia left her home and began walking at 6 km/hr toward Kiley’s house. Fifteen minutes later, Kiley left her home and walked at 9 km/hr toward Lucia’s house. If Lucia lives 5 km from Kiley, at what time did they meet? Who walked farther?

Answer by jorel1380(3719)

6k per hour x 1/4 hour=1.5 k's
5-1.5=3.5
9n+6n=3.5
15n=3.5
n=.2333 hours
Lucia walked 1.5+.233333(6)=2.9 k's
Kiley walked .2333333(9)=2.1 k's
Lucia walked further..